# Geometric Scaling of Bayesian Inference in LLMs

### v2

@@ 4.6. Causal Intervention Protocol: Entropy-Axis Ablations
For the causal probes in Section Section 6, we construct entropy-aligned axes and ablate them using a simple linear projection scheme.
@@
For each layer $\ell$ of Pythia-410M we collect last-layer value vectors $v^{(\ell)}_{T,h}$ for the SULA prompts and compute a layer-specific PCA basis as in Section Section 4.2. We take the first principal component $u^{(\ell)}_{\text{ent}}$ and orient its sign so that the dot product $v^{(\ell)}_{T,h}\cdot u^{(\ell)}_{\text{ent}}$ is negatively correlated with predictive entropy; this defines the entropy axis for that layer. Random control axes $u^{(\ell)}_{\text{rand}}$ are drawn by sampling a Gaussian vector in the same space and normalizing to unit norm.
@@ Results.
Table Table 1 and Figure Figure 1 summarize the findings. The domain-restriction effect is model-dependent rather than universal:
@@
  • Pythia-410M exhibits near-complete dimensionality collapse under both conditions (PC1+PC2 $\approx$ 100%), with no significant difference between mixed and domain-restricted prompts. This suggests that Pythia’s value manifold operates in a consistently low-dimensional subspace regardless of domain.
  • Phi-2 shows moderate dimensionality (PC1+PC2 $\approx$ 60–64%) with minimal domain effect. The curated training data may induce stable geometric structure that is robust to prompt domain.
@@
The wind-tunnel experiments (Paper 1) used tasks with a single, analytically tractable posterior—effectively a maximally domain-restricted setting. The Llama results demonstrate that this pattern can emerge in production models when prompt distributions are similarly constrained. However, the Pythia and Phi-2 results show that not all architectures exhibit this behavior, suggesting that the mapping from domain restriction to manifold collapse depends on training dynamics and architectural capacity.
@@ 5.2. Inference-Time Bayesian Updating in Production Models (SULA)
Each SULA prompt contains $k$ labeled examples of the form $x_{i}$ is positive” or $x_{i}$ is negative”, followed by a query word $x_{\mathrm{query}}$. Labels carry no semantic content; they serve only as discrete likelihood indicators. We use a simple binary hypothesis model with likelihood ratio 0.9:0.10.9{:}0.1. Let $y\in\{\text{positive},\text{negative}\}$ denote the latent class. The prior is uniform, and each example contributes independent evidence:
$p(y\mid\mathcal{D}_{k})\;\propto\;p_{0}(y)\prod_{i=1}^{k}\ell(x_{i},y_{i}),\qquad\ell(x_{i},y_{i})=\begin{cases}0.9&\text{if }y=y_{i},\\[3.0pt] 0.1&\text{if }y\neq y_{i}.\end{cases}$
The resulting analytical posterior entropy $H_{\mathrm{Bayes}}(k)$ is known in closed form and decreases monotonically with $k$.
@@
We generate 250 prompts for each $k\in\{0,1,2,4,8\}$ with varying label imbalances. For each prompt, we extract: (1) the model’s predictive entropy, (2) last-layer value vectors projected into a common PCA basis (Section Section 4.2), and (3) attention-entropy trajectories. This allows us to test whether value-manifold coordinates move along the Bayesian axis as evidence accumulates.
@@
Figure Figure 2 summarizes the findings.
Predictive entropy. Model entropy declines monotonically with $k$ and tracks analytical Bayesian entropy:
@@
Although noisier than wind-tunnel calibration, the consistent monotone trend indicates that production models extract and use likelihood information supplied in the prompt.
Manifold alignment. When all SULA value vectors (across all $k$) are embedded into a shared PCA basis, PC1 coordinates correlate strongly with the analytical Bayesian entropy for each model ($|\rho|=0.65$—-$0.80$). This confirms that the entropy-ordered manifold learned during large-scale training is the axis along which inference-time updates occur.
Bayesian-axis trajectory. The mean PC1 coordinate shifts monotonically with $k$:
@@
Control conditions (monotone SULA).
To verify that manifold movement reflects likelihood structure rather than prompt format, we implement three control conditions tailored to the monotone SULA generative model. In all conditions, the latent label $y$ and the number of glimpses $k$ are preserved so that analytic Bayesian entropies remain well-defined.
@@
    Lexical remapping. We preserve the latent label $y$ and the full 0.9:0.1 noisy-glimpses generative model but replace the two surface label tokens $(L_{1},L_{2})$ with a different fixed pair $(L_{1}^{\prime},L_{2}^{\prime})$. This tests whether manifold trajectories depend on the identity of the nonsense tokens rather than on the underlying likelihood information. Analytical posteriors are unchanged.
@@
    Shuffled labels. We preserve the prompt template but overwrite each example’s surface label token with an independent draw from $\{L_{1},L_{2}\}$, breaking all correlation between glimpses and the latent label. Under the SULA generative model, this corresponds to providing no usable evidence, so the analytical posterior collapses to the uniform prior for all $k$ ($H_{\mathrm{Bayes}}=1$ bit). Any systematic manifold movement should therefore disappear.
@@
    Evidence ablation. We retain the surface labels and prompt structure but mask the glimpse content (e.g., replacing carrier tokens with “[MASK]”). This removes the likelihood-bearing evidence while preserving superficial formatting, again yielding a flat analytical posterior. This tests whether models move along the manifold only when evidence tokens encode useful likelihood information.
@@
The SULA experiment demonstrates that production models use the same geometric mechanism active in wind-tunnel transformers: evidence supplied in-context drives representations along an entropy-ordered manifold. The calibration gap (0.31–0.44 bits vs. <0.1<0.1 bits in wind tunnels) reflects the fact that (i) production models are not trained on the SULA generative distribution, and (ii) natural-language prompts introduce semantic ambiguity absent in synthetic tasks. The key result is the systematic correspondence between analytical Bayesian entropy, model predictive entropy, and movement along the value manifold.
Together, these findings show that geometric Bayesian updating is an inference-time phenomenon: transformers navigate the same manifold direction that encodes predictive uncertainty when supplied with usable likelihood information inside the prompt.
@@ 5.6. Scaling Within a Family: Pythia-12B
Mixed-domain geometry becomes multi-lobed (PC1+PC2 = 19%), but mathematics-only prompts recover a near-1D manifold (PC1+PC2 $\approx$ 0.90).
@@ 5.7. Boundary Case: The Mistral Family
All Mistral variants exhibit low-dimensional value manifolds under mixed-domain prompts and recover wind-tunnel-style 1D collapse under mathematics-only prompts ($_{1}{+}PC$$_{2}\approx 80\%$$90\%$). Key orthogonality is likewise sharp: early and mid layers show mean off-diagonal cosine values near 0.050.050.060.06, well below both Gaussian and initialization baselines. These results indicate that the hypothesis-frame structure and entropy-ordered manifold discovered in Papers 1–2 persist robustly in Mistral architectures.
@@
Figure 7. Attenuated dynamic focusing in Mistral-style architectures. Attention entropy as a function of layer depth for three variants of the Mistral family. Unlike models with full-sequence multi-head attention, entropy decreases only modestly ($20\%$$30\%$) and often non-monotonically, reflecting weakened dynamic routing due to (i) sliding-window attention, which prevents global evidence accumulation, and (ii) mixture-of-experts routing, which fragments updates across experts. Despite reduced focusing dynamics, the static value-space geometry (Sections Section 5.1Section 5.2) remains intact, illustrating a dissociation between representational invariants and the mechanisms that refine them during inference.
@@
This pattern matches the frame–precision dissociation predicted in Paper 2: attention patterns (the frame) stabilize early and robustly under training, whereas the precision of posterior refinement is sensitive to architectural bandwidth and routing design.
@@ Conclusion.
The Mistral family should therefore be viewed as a boundary case that reveals how architectural constraints selectively modulate dynamic Bayesian computation, rather than as a counterexample to Bayesian geometry. Static geometric structure—entropy-ordered manifolds and hypothesis frames—persists across all Mistral variants, while dynamic refinement is diminished by local attention and MoE routing. This provides a natural explanation for the observed behavior and connects directly to the theoretical predictions of Paper 2.
@@ 5.8. Cross-Architecture Synthesis
Table Table 2 summarizes the unified picture:
@@
  • Attention focusing: layerwise entropy reduction varies systematically by architecture—-strong in full-sequence MHA, moderate in GQA, and modest and often non-monotone in sliding-window and MoE variants.
As summarized in Table Table 2, all three geometric signatures show quantitative continuity from synthetic wind-tunnel tasks to production models.
@@
The stable invariants across every architecture - including Mistral - are: (1) entropy-ordered value manifolds, and (2) orthogonal key frames. Dynamic focusing is architecture-dependent, requiring global routing capacity.
This separation is the central representational—-computational split predicted in Paper 2.
Figure 8. Cross-model geometric signatures. Normalized comparison of four geometric metrics across three model families. Left: Static signatures—-value manifold collapse ($_{1}{+}PC$2) and domain-restriction gain—-are consistently present, indicating that all models internalize a low-dimensional representation of hypothesis space. Right: Dynamic signatures—-key orthogonality and attention focusing—-vary substantially by architecture: Llama-3.2-1B exhibits strong runtime refinement, whereas Pythia-410M relies on pre-orthogonalized hypothesis frames and shows weak focusing. This dissociation suggests that representation is universal, but the mechanism that refines it during inference is architectural.
@@ 6.1. Universal core mechanisms
Across the standard MHA and GQA models (Pythia, Phi-2, and Llama), we observe the full set of Bayesian geometric signatures predicted by the wind tunnel experiments: value manifolds, key orthogonality, and layerwise attention focusing. In the Mistral family (Section Section 5.7), the static signatures (value manifolds and key orthogonality) persist, while the dynamic signature (monotone attention focusing) is weakened or absent due to architectural and training-objective differences. This suggests a natural split between universal static structure and architecture-dependent dynamics.
Value manifolds (static). All models, including the Mistral variants, show low-dimensional value structure, with PC1+PC2 between roughly 16% and 84%, far above random baselines. Entropy ordering persists in every model evaluated: prompts with higher next-token entropy occupy systematically different regions of the manifold than low-entropy prompts.
Key orthogonality (static). All models learn hypothesis-frame structure, with mean off-diagonal cosine between 0.034 and 0.18, consistently 2 - 10 times better than random initialization (0.40 - 0.45). Mistral models show the same early-layer emergence and late-layer collapse of orthogonality as standard MHA.
Attention focusing (dynamic). Standard MHA and GQA models exhibit a clear layerwise entropy decrease (31 - 86%), matching the binding $\rightarrow$ elimination $\rightarrow$ refinement pattern seen in wind tunnels. In the Mistral models, this dynamic signature is weak or absent (Section Section 5.7), consistent with architectural constraints (sliding-window attention, MoE routing) and the effects of downstream alignment.
The ICL experiment connects the controlled wind-tunnel setting to real-time inference in production models. In the wind tunnels, Bayesian structure emerges from the training objective alone: the model learns to encode posterior uncertainty along a one-dimensional manifold. The ICL setting demonstrates that the same mechanism is active during inference: when the model is given explicit symbolic evidence inside the prompt, its value representations move along the same manifold direction that encodes posterior entropy, and the degree of movement is proportional to the amount of evidence. In this way, the ICL results show that transformers not only embody Bayesian geometry as a representational primitive, but also execute Bayesian updates on-the-fly when the prompt supplies usable likelihood information.
Conclusion. The static components of the Bayesian geometry (value manifolds and hypothesis frames) are universal across architectures, while strong progressive attention focusing depends on global routing capacity and training regime.
@@ 6.2. Training Data Quality Enhances Clarity
High-quality, consistent training examples enable gradient dynamics to sculpt sharper hypothesis frames (better orthogonality) and stronger attention pathways (better focusing). This suggests:
Implication for training: Early training on curated data may establish cleaner geometric scaffolding that persists through subsequent web-scale training. Curriculum learning - progressing from structured to diverse data - could enhance both interpretability and reliability.
@@
Our final experiment simultaneously ablates the entropy-aligned value direction across all identified “Bayesian layers” (8, 12, 16, 20, 23) in Pythia-410M. If this axis encoded a causal bottleneck for Bayesian inference, removing it at every layer should degrade and ultimately collapse SULA behavior. Instead, we observe the opposite dissociation: (i) the multi-layer ablation destroys the entropy geometry (axis—-entropy correlation drops from $0.27$ to $0.07$), yet (ii) SULA calibration remains intact, with MAE and correlation to the Bayesian posterior changing by less than $1\%$, and (iii) a matched random-axis ablation improves calibration metrics despite preserving the original geometry. These results rule out both single-axis and multi-axis bottleneck hypotheses and show that the entropy manifold is a representational trace of inference rather than the substrate that performs it.

### v3

@@ 4.6. Causal Intervention Protocol: Entropy-Axis Ablations
For the causal probes in Section 6, we construct entropy-aligned axes and ablate them using a simple linear projection scheme.
@@
For each layer $\ell$ of Pythia-410M we collect last-layer value vectors $v^{(\ell)}_{T,h}$ for a held-out set of 200 SULA prompts (disjoint from the 300 evaluation prompts) and compute a layer-specific PCA basis as in Section 4.2. We take the first principal component $u^{(\ell)}_{\text{ent}}$ and orient its sign so that the dot product $v^{(\ell)}_{T,h}\cdot u^{(\ell)}_{\text{ent}}$ correlates positively with predictive entropy (i.e., high projection = high uncertainty); this defines the entropy axis for that layer. Interventions are then evaluated on the separate evaluation set to avoid overfitting. Random control axes $u^{(\ell)}_{\text{rand}}$ are drawn by sampling a Gaussian vector in the same space and normalizing to unit norm.
@@ Results.
Table 1 and Figure 1 summarize the findings. The domain-restriction effect is model-dependent rather than universal:
@@
  • Pythia-410M exhibits near-complete dimensionality collapse under both conditions (PC1+PC2 $\approx$ 100%), with no significant difference between mixed and domain-restricted prompts. Its value manifold operates in a consistently low-dimensional subspace regardless of domain.
  • Phi-2 shows moderate dimensionality (PC1+PC2 $\approx$ 60–64%) with minimal domain effect. The curated training data may induce stable geometric structure that is stable across prompt domains.
@@
The wind-tunnel experiments (Paper I) used tasks with a single, analytically tractable posterior—effectively a maximally domain-restricted setting. The Llama results demonstrate that this pattern can emerge in production models when prompt distributions are similarly constrained. However, the Pythia and Phi-2 results show that not all architectures exhibit this behavior, suggesting that the mapping from domain restriction to manifold collapse depends on training dynamics and architectural capacity.
@@ 5.2. Inference-Time Bayesian Updating in Production Models (SULA)
Each SULA prompt contains $k$ labeled examples of the form $x_{i}$ is positive” or $x_{i}$ is negative”, followed by a query word $x_{\mathrm{query}}$. Labels carry no semantic content; they serve only as discrete likelihood indicators. We compute analytical Bayesian posteriors using a Beta-Bernoulli model. Let $\theta\in[0,1]$ be the latent probability that a word is positive. We place a uniform Beta(1,1) prior on $\theta$. For each labeled example, the likelihood is:
[…]
The posterior $p(\theta\mid k\text{ examples})$ is computed via Bayes’ rule, and we report its entropy $H_{\mathrm{Bayes}}(k)$, which is known in closed form and decreases monotonically with $k$. Sentiment words are drawn from a vocabulary of 50 positive and 50 negative words; labels are sampled with 70% consistency (matching true sentiment) and 30% inconsistency.
@@
We generate 250 prompts for each $k\in\{0,1,2,4,8\}$ with varying label imbalances. For each prompt, we extract: (1) the model’s predictive entropy, (2) last-layer value vectors projected into a common PCA basis (Section 4.2), and (3) attention-entropy trajectories. This allows us to test whether value-manifold coordinates move along the Bayesian axis as evidence accumulates.
@@
Figure 2 summarizes the findings. Predictive entropy. Model entropy declines monotonically with $k$ and tracks analytical Bayesian entropy:
@@
Although noisier than wind-tunnel calibration, the consistent monotone trend indicates that production models extract and use likelihood information supplied in the prompt. Manifold alignment. When all SULA value vectors (across all $k$) are embedded into a shared PCA basis, PC1 coordinates correlate strongly with the analytical Bayesian entropy for each model ($|\rho|=0.65$$0.80$). This confirms that the entropy-ordered manifold learned during large-scale training is the axis along which inference-time updates occur. Bayesian-axis trajectory. The mean PC1 coordinate shifts monotonically with $k$:
@@
Control conditions (SULA).
To verify that manifold movement reflects likelihood structure rather than prompt format, we implement three control conditions. All controls should eliminate the correlation between manifold position and Bayesian entropy.
@@
    Lexical remap. Replace sentiment words with random tokens, breaking semantic meaning while preserving syntax. This tests whether manifold trajectories depend on the identity of the tokens rather than on the underlying likelihood information. Analytical posteriors are unchanged.
@@
    Shuffled labels. Randomly permute labels across examples, breaking the label-word correspondence. Under the SULA generative model, this corresponds to providing no usable evidence, so the analytical posterior collapses to the uniform prior for all $k$. Any systematic manifold movement should therefore disappear.
@@
    Evidence ablation. Mask the labeled examples entirely, leaving only the query. This removes the likelihood-bearing evidence while preserving superficial formatting, again yielding a flat analytical posterior. This tests whether models move along the manifold only when evidence tokens encode useful likelihood information.
@@
The SULA experiment demonstrates that production models use the same geometric mechanism active in wind-tunnel transformers: evidence supplied in-context drives representations along an entropy-ordered manifold. The calibration gap (0.31–0.44 bits vs. <0.1<0.1 bits in wind tunnels) reflects the fact that (i) production models are not trained on the SULA generative distribution, and (ii) natural-language prompts introduce semantic ambiguity absent in synthetic tasks. The key result is the systematic correspondence between analytical Bayesian entropy, model predictive entropy, and movement along the value manifold. Together, these findings show that geometric Bayesian updating is an inference-time phenomenon: transformers navigate the same manifold direction that encodes predictive uncertainty when supplied with usable likelihood information inside the prompt.
@@ 5.6. Scaling Within a Family: Pythia-12B
Mixed-domain geometry becomes multi-lobed (PC1+PC2 = 19%), but mathematics-only prompts recover a near-1D manifold (PC1+PC2 $\approx$ 0.90). Pythia-12B shows markedly different mixed-domain geometry than Pythia-410M (19% vs 99.7%), suggesting scale increases representational distribution. Under domain restriction both converge to similar collapse ($\sim$}90%), indicating the geometric substrate is preserved but more distributed at scale.
@@ 5.7. Boundary Case: The Mistral Family
All Mistral variants exhibit low-dimensional value manifolds under mixed-domain prompts and recover wind-tunnel-style 1D collapse under mathematics-only prompts ($_{1}{+}PC$$_{2}\approx 80\%$$90\%$). Key orthogonality is likewise sharp: early and mid layers show mean off-diagonal cosine values near 0.050.050.060.06, well below both Gaussian and initialization baselines. These results indicate that the hypothesis-frame structure and entropy-ordered manifold discovered in Papers I–II persist in Mistral architectures.
@@
Figure 7. Attenuated dynamic focusing in Mistral-style architectures. Attention entropy as a function of layer depth for three variants of the Mistral family. Unlike models with full-sequence multi-head attention, entropy decreases only modestly ($20\%$$30\%$) and often non-monotonically, reflecting weakened dynamic routing due to (i) sliding-window attention, which prevents global evidence accumulation, and (ii) mixture-of-experts routing, which fragments updates across experts. Despite reduced focusing dynamics, the static value-space geometry (Sections 5.1 and 5.2) remains intact, illustrating a dissociation between representational invariants and the mechanisms that refine them during inference.
@@
This pattern matches the frame–precision dissociation predicted in Paper II: attention patterns (the frame) stabilize early and robustly under training, whereas the precision of posterior refinement is sensitive to architectural bandwidth and routing design.
@@ Conclusion.
The Mistral family should therefore be viewed as a boundary case that reveals how architectural constraints selectively modulate dynamic Bayesian computation, rather than as a counterexample to Bayesian geometry. Static geometric structure—entropy-ordered manifolds and hypothesis frames—persists across all Mistral variants, while dynamic refinement is diminished by local (sliding-window) attention. This provides a natural explanation for the observed behavior and connects directly to the theoretical predictions of Paper II.
@@ 5.8. Cross-Architecture Synthesis
Table 2 summarizes the unified picture:
@@
  • Attention focusing: layerwise entropy reduction varies systematically by architecture–strong in full-sequence MHA, moderate in GQA, and modest and often non-monotone in sliding-window and MoE variants.
As summarized in Table 2, all three geometric signatures show quantitative continuity from synthetic wind-tunnel tasks to production models.
@@
The stable invariants across every architecture - including Mistral - are: (1) entropy-ordered value manifolds, and (2) orthogonal key frames. Dynamic focusing is architecture-dependent, requiring global routing capacity. This separation is the central representational–computational split predicted in Paper II.
Figure 8. Cross-model geometric signatures. Normalized comparison of four geometric metrics across three model families. Left: Static signatures—value manifold collapse ($_{1}{+}PC$2) and domain-restriction gain—are consistently present, indicating that all models internalize a low-dimensional representation of hypothesis space. Right: Dynamic signatures—key orthogonality and attention focusing—vary substantially by architecture: Llama-3.2-1B exhibits strong runtime refinement, whereas Pythia-410M relies on pre-orthogonalized hypothesis frames and shows weak focusing. This dissociation suggests that representation is universal, but the mechanism that refines it during inference is architectural.
@@ 6.1. Universal core mechanisms
Across the standard MHA and GQA models (Pythia, Phi-2, and Llama), we observe the full set of Bayesian geometric signatures predicted by the wind tunnel experiments: value manifolds, key orthogonality, and layerwise attention focusing. In the Mistral family (Section 5.7), the static signatures (value manifolds and key orthogonality) persist, while the dynamic signature (monotone attention focusing) is weakened or absent due to architectural and training-objective differences. This suggests a natural split between universal static structure and architecture-dependent dynamics. Value manifolds (static). All models, including the Mistral variants, show low-dimensional value structure, with PC1+PC2 between roughly 16% and 84%, far above random baselines. Entropy ordering persists in every model evaluated: prompts with higher next-token entropy occupy systematically different regions of the manifold than low-entropy prompts. Key orthogonality (static). All models learn hypothesis-frame structure, with mean off-diagonal cosine between 0.034 and 0.18, consistently 2 - 10 times better than random initialization (0.40 - 0.45). Mistral models show the same early-layer emergence and late-layer collapse of orthogonality as standard MHA. Attention focusing (dynamic). Standard MHA and GQA models exhibit a clear layerwise entropy decrease (31–86%), matching the binding $\rightarrow$ elimination $\rightarrow$ refinement pattern seen in wind tunnels. In the Mistral models, this dynamic signature is weak or absent (Section 5.7), consistent with architectural constraints (sliding-window attention limiting global context access) and the effects of downstream alignment. The 50+ percentage point gap between MHA models (82–86%) and GQA/sliding-window models (20–31%) indicates that architectural constraints can substantially limit dynamic Bayesian refinement even when static geometry is preserved. The ICL experiment connects the controlled wind-tunnel setting to real-time inference in production models. In the wind tunnels, Bayesian structure emerges from the training objective alone: the model learns to encode posterior uncertainty along a one-dimensional manifold. The ICL setting demonstrates that the same mechanism is active during inference: when the model is given explicit symbolic evidence inside the prompt, its value representations move along the same manifold direction that encodes posterior entropy, and the degree of movement is proportional to the amount of evidence. In this way, the ICL results show that transformers not only embody geometry consistent with Bayesian inference as a representational primitive, but also execute updates correlated with Bayesian posteriors when the prompt supplies usable likelihood information. Conclusion. The static components of the Bayesian geometry (value manifolds and hypothesis frames) are universal across architectures, while strong progressive attention focusing depends on global routing capacity and training regime.
@@ 6.2. Training Data Quality Enhances Clarity
High-quality, consistent training examples enable gradient dynamics to sculpt sharper hypothesis frames (better orthogonality) and stronger attention pathways (better focusing). Implication for training: Early training on curated data may establish cleaner geometric scaffolding that persists through subsequent web-scale training. Curriculum learning - progressing from structured to diverse data - could enhance both interpretability and reliability.
@@
Our final experiment simultaneously ablates the entropy-aligned value direction across all identified “Bayesian layers” (8, 12, 16, 20, 23) in Pythia-410M. If this axis encoded a causal bottleneck for Bayesian inference, removing it at every layer should degrade and ultimately collapse SULA behavior. Instead, we observe the opposite dissociation: (i) the multi-layer ablation destroys the entropy geometry (axis–entropy correlation drops from $0.27$ to $0.07$), yet (ii) SULA calibration remains intact, with MAE and correlation to the Bayesian posterior changing by less than $1\%$, and (iii) a matched random-axis ablation improves calibration metrics despite preserving the original geometry. These results rule out both single-axis and multi-axis bottleneck hypotheses and show that the entropy manifold is a representational trace of inference rather than the substrate that performs it.